The Forsite Company is screening three ideas for new services. Resource constraints allow only one idea to be commercialized at the present time. The following estimates have been made for the five performance criteria that management believes to be most​ important:

                                                                                                                                                                                                                                                  

Rating

Performance Criterion

Service A

Service B

Service C

Capital equipment investment required

0.30.3

0.70.7

0.60.6

Expected return on investment​ (ROI)

0.60.6

0.50.5

0.20.2

Compatibility with current workforce skills

0.40.4

0.60.6

0.50.5

Competitive advantage

0.60.6

0.70.7

0.70.7

Compatibility with EPA requirements

0.90.9

0.80.8

0.20.2

a. Calculate a total weighted score for each alternative. Use a preference matrix and assume equal weights for each performance criterion. Let the total of the weights equal 100.

​(Enter

your responses as

integers.​)
Answer #1

Service A = 20*0.3 + 20*0.6 + 20*0.4 + 20*0.6 + 20*0.9

= (6+12+8+12+18)

= 56

Service B = 20*0.7 + 20*0.5 + 20*0.6 + 20*0.7 + 20*0.8

= (14+10+12+14+16)

= 66

Service C = 20*0.6 + 20*0.2 + 20*0.5 + 20*0.7 + 20*0.2

= (12 + 4 + 10 + 14 + 4)

= 44